OpenStudy (anonymous):

Write the equation of the line that passes through (-4, 12) and (7, 23). A. y = x + 16 B. y = 3x + 18 C. y = 1/6 x + 16 D. y = 1/3 x + 18

OpenStudy (anonymous):

m=(23-12)/(7+4)=1 y-12=1(x+4) y=x+16 So it was A Hope this helps ! :)